OpenStudy (amistre64):

abstract algebra: what is the difference between an "orbit" and a "cycle" regarding permutations?

OpenStudy (abb0t):

cycle goes on forever, like trig functions. One cycle is from 0 to 2pi, but orbit continuously goes around same path?

OpenStudy (amistre64):

is an orbit a collection of cycles that define a single permuation? all orbits are cycles. a cycle is a set of directions, and can suggest more than a one step permutation?

OpenStudy (amistre64):

\[\sigma=\begin{pmatrix}1&2&3&4&5\\3&5&4&1&2\end{pmatrix}=(134)(25)\]

OpenStudy (amistre64):

the orbits are: (134) and (25); or is that {1,3,4} and {2,5} proper notation?

OpenStudy (amistre64):

product of 3 cycles \[\mu=(126)(24)(31)=\begin{pmatrix} 1&2&3&4&5&6\\ 3&2&1&4&5&6\\ 3&4&1&2&5&6\\ 3&4&2&6&5&1\\ \end{pmatrix}= \begin{pmatrix} 1&2&3&4&5&6\\ 3&4&2&6&5&1\\ \end{pmatrix}\]

OpenStudy (anonymous):

i think the difference is that orbits are sets the ideas are the same

OpenStudy (anonymous):

also an "orbit" can be on element, whereas i don't think you count (a) as a cycle

OpenStudy (anonymous):

*one element

OpenStudy (amistre64):

is the proper notation for an orbit a roster set notation then? {a,b,c,...,n}

OpenStudy (amistre64):

also, product of cycles, if seen them go left to right, and other right to left ... not sure if there is a proper direction for those